Two equal forces, 78.6 N each, gives you a resultant value of 78.6 N. What is the angle between the two forces?
Answer #1

78.6 = sqroot ( 78.6^2 + 78.6^2 - 2 ( 78.6)^2 cos theta)

78.6^2 = 2 ( 78.6^2) - 2 ( 78,6^2) cos theta

78,6^2 = 2( 78.6^2) cos theta

angle = 60